OpenStudy (anonymous):

Solve the system algebraically. 2x + y - 10 = 0 x - y - 5 = 0 What is the value of y? -5/3 0 5

OpenStudy (anonymous):

i really need help

OpenStudy (anonymous):

i can help! :)

OpenStudy (anonymous):

great

OpenStudy (anonymous):

y=0 but i will explain why

OpenStudy (anonymous):

\[2 x + 1 y = 10 \]

OpenStudy (anonymous):

\[1 x -1 y = 5 \]

OpenStudy (anonymous):

Eliminate y multiply both by 1

OpenStudy (anonymous):

\[2 x + 1 y = 10 \]

OpenStudy (anonymous):

\[1 x + -1 y = 5 \]

OpenStudy (anonymous):

Add the two equations

OpenStudy (anonymous):

\[3x=\frac{ 15 }{ 3 }\]

OpenStudy (anonymous):

\[x = 5 \]

OpenStudy (anonymous):

plug value of x in \[2 x + 1 y = 10 \]

OpenStudy (anonymous):

\[10 + 1 y = 10 \]

OpenStudy (anonymous):

\[1 y = 10 -10\]

OpenStudy (anonymous):

\[y = 0 \]